OpenStudy (anonymous):

\[Find~ the~ value~ of ~~\alpha~~ btween ~{1\over2}\pi\]\[for~which ~~3sinx+2cosx=\sqrt{13}\sin(x+\alpha)\]

OpenStudy (anonymous):

\[btween~~ 0 ~and ~{1\over2} \pi\]

OpenStudy (anonymous):

\[\tan \alpha=\frac{2}{3}\]

OpenStudy (anonymous):

I don't think that's the right answer?

OpenStudy (anonymous):

It says 0.588 is the answer...

OpenStudy (mimi_x3):

tan inverse of 2/3 is 0.588..

OpenStudy (anonymous):

Yeah it is inverse of tan in radian measure

OpenStudy (anonymous):

Oh, yes. I was working in degrees.

OpenStudy (anonymous):

How did you get it?

OpenStudy (mimi_x3):

I used a calculator, lol \[a = \tan^{-1} (\frac{2}{3} )\]

OpenStudy (anonymous):

No, lol. I mean tanα=2/3?

OpenStudy (mimi_x3):

Auxilliary Angle.. \[asinx + bcosx = \sqrt{a^2+b^2} \sin(x+\alpha) \] \[\tan \alpha = \frac{b}{a} \]

OpenStudy (anonymous):

Multiply by sqrt(13) NOW you have 2/sqrt(13) and 3/sqrt(13) beside sinx and cosx these are sides of a triangle  you get some insight

OpenStudy (anonymous):

@Mimi_x3 I don't completely understand you. @NotSObright My visual thinking is very bad... (in other words~ I can't work with diagrams) lol

OpenStudy (mimi_x3):

What don't you understand ? It's the formula..

OpenStudy (anonymous):

I understand the formula... but I don't know how you arrive at tanα=b/a with the formula...

OpenStudy (anonymous):

\[\cos(\alpha)=\frac{2}{\sqrt13}\] \[\sin(\alpha)=\frac{3}{\sqrt13}\] subtitute this to get the compund angle formula

OpenStudy (anonymous):

I can see how tanα=b/a becomes tanα=2/3 and then the answer... but not how asinx+bcosx=√a^22+b^22sin(x+α) ~~~~~~~~~~~~ tanα=b/a?

OpenStudy (campbell_st):

given a sin (x) + b cos(x) it can be rewritten as \[r \sin(x + \alpha)\] \[r = \sqrt{a^2 + b^2}\] pythagorean identity \[\tan \alpha = b/a\] in your question \[r = \sqrt{3^2 + 2^2} \] \[r = \sqrt{13}\] \[\tan \alpha = 2/3\] \[\alpha = \arctan (2/3)\] \[\alpha = 0.588 \] radians then \[\sqrt{13} \sin(x + 0.588)\]

OpenStudy (anonymous):

Ok, but how did you get tanα=b/a?